# Random Stability of Random Variables

### 30 Mar 2026 v1

@@ 7.1. Proof of Theorem 1
It is real-valued, since $\Sigma_{k}\stackrel{{\scriptstyle d}}{{=}}-\Sigma_{k}$. It turns out that this simplifies the proof in a critical way. See [38, Chapter 3], Theorem 3.1.2. Applying [54, Chapter 3], Lemma 3.1.3, we derive from tightness of the sequence $(\Sigma_{k})$ that
@@
Since $g_{k}$ is real-valued, and by [38, Chapter 2], page 36, $g_{k}(u)\leq 1$, we can simply write this statement (39) without the absolute value. This characteristic function $g_{k}$ from (38) can be represented as
@@
is the characteristic function of $Y_{n}$, see [38, Chapter 3], Corollary 2 of Theorem 3.3.1.
Step 5. It follows from scaling theory for discrete-time branching processes, see for example [50, Theorem 4.4] with infinite mean that for a fixed $a>1$, the event $A=\{N_{k}/a^{k}\to\infty\}$ has positive probability. For any $s\in(0,1)$, we can rewrite
@@
In the right-hand side of (44), the number $s\in(0,1)$ is arbitrary. Therefore, we can take $s$ as close to $1$ as we wish, to make $\ln s$ negative but as close to zero as we wish. Also, recall that by [38, Chapter 2], page 36, we have: $g(uc^{-k})\leq 1$. Applying all this to (44), we get:
@@
where we define $\triangle^{t}_{2}g(u):=g(u-2t)+g(u+2t)-2g(u)$. Clearly, $t_{k}\to 0$, and letting $a=c^{2}$ (since $a$ is arbitrary), we rewrite (45) as $\triangle_{2}^{t_{k}}g(0)/t_{k}^{2}\to 0$. Apply [38, Chapter 2], Theorem 2.3.1 and conclude: $g^{\prime\prime}(0)=0$, thus $\mathbb{E}[Y^{2}]=0$, and $Y=0$ almost surely, thus $g(u)\equiv 1$. Comparing this with (40), we get: $|f(u)|\equiv 1$. Apply [38, Chapter 2], Theorem 2.1.4, Corollary 1. We get: $f(u)=e^{\mathrm{i}uX_{0}}$ and $X=x_{0}$ almost surely for some $x_{0}\in\mathbb{R}$. But $X$ is $N$-stable. Plugging $X=x_{0}$ into (1), we get: $x_{0}N=cx$. Thus either $N=c$, which contradicts $\mathbb{E}[N]=\infty$; or $x_{0}=0$, which implies $X=0$ almost surely, but we exclude this trivial case. This completes the proof that there is no $N$-stable $X$ in case $\mathbb{E}[N]=\infty$.
@@ 7.2. Proof of Theorem 2
for some $g\in\mathcal{G}$ of index $\alpha$ can be represented as $X=Y^{1/\alpha}Z$, where $Y$ is the standard $N$-stable random variable, and $\mathbb{E}\left[e^{\mathrm{i}uZ}\right]=e^{-g(u)}$. We simply apply [32, Proposition 3.1] and use their remarks on page 309, (10) – (12), Remark 1, about the difference between stable and strictly stable. One can also consult [47, Chapter 1], Section 2 for the latter question.
@@ 7.3. Proof of Theorem 3
If $X$ is a constant, then $M$ and $N$ are also constants, and obviously they commute. If $X$ is not a constant, then $f$ takes at least one value $z$ inside the unit disc $\mathbb{D}=\{z\in\mathbb{C}\mid|z|<1\}$, [38, Chapter 2], Theorem 2.1.4, Corollary 2. But $f$ is continuous, therefore this value $z$ is the limit point of the image of $f$. Both $\varphi\circ\psi$ and $\psi\circ\varphi$ are probability generating functions of some distributions. Therefore, they are analytic on $\mathbb{D}$. Applying the classic result from compelx analysis, we get (16).
@@ 7.7. Proof of Lemma 10
Since $X\geq 0$, by assumption, then the strictly stable random variable $Z$ with index $\alpha$ must be nonnegative as well. But this could be true only if $\alpha=1$ and $Z$ is a positive constant; or $\alpha\in(0,1)$, and $Z$ has Laplace transform (10). This can be found in [53, Chapter V], Theorem 3.5, or in the classic reference [47]. Then the Laplace transform of $X$ is
@@ 8. Conclusions and Open Problems
Further possible research directions include generalizing the results for many dimensions. The most obvious generalization is simply to rewrite the main equation (1) with vector $X$ but scalar $N$ and $c$. But results are transferred verbatim to this case; see [22, Chapter 10], [26] for multivariate geometric stable distributions. Thus we chose not to write a separate section on this topic. However, there exists a more substantial generalization, pioneered by T. Kozubowski and A. Panorska: operator stable laws, when the $c$ from (1) is a constant matrix, not a scalar. See [30] for operator geometric stable laws, and [31] for general operator random stable laws. We propose to extend our results to operator random stable distributions.
@@ Acknowledgements
We thank our departmental colleague Tomasz Kozubowski for raising the question whether there exists a Sibuya-stable random variable, and another question whether there exist non-geometric and non-Sibuya commuting semigroups; and further useful discussion. We thank Thierry Huillet for pointing out recent articles on continuous-time branching processes with explicit distributions. We thanks Svetlozar Rachev for multiple useful comments on an earlier draft.

### 02 Apr 2026 v2

@@ 7.1. Proof of Theorem 1
It is real-valued, since $\Sigma_{k}\stackrel{{\scriptstyle d}}{{=}}-\Sigma_{k}$. It turns out that this simplifies the proof in a critical way. See [36, Chapter 3], Theorem 3.1.2. Applying [52, Chapter 3], Lemma 3.1.3, we derive from tightness of the sequence $(\Sigma_{k})$ that
@@
Since $g_{k}$ is real-valued, and by [36, Chapter 2], page 36, $g_{k}(u)\leq 1$, we can simply write this statement (39) without the absolute value. This characteristic function $g_{k}$ from (38) can be represented as
@@
is the characteristic function of $Y_{n}$, see [36, Chapter 3], Corollary 2 of Theorem 3.3.1.
Step 5. It follows from scaling theory for discrete-time branching processes, see for example [48, Theorem 4.4] with infinite mean that for a fixed $a>1$, the event $A=\{N_{k}/a^{k}\to\infty\}$ has positive probability. For any $s\in(0,1)$, we can rewrite
@@
In the right-hand side of (44), the number $s\in(0,1)$ is arbitrary. Therefore, we can take $s$ as close to $1$ as we wish, to make $\ln s$ negative but as close to zero as we wish. Also, recall that by [36, Chapter 2], page 36, we have: $g(uc^{-k})\leq 1$. Applying all this to (44), we get:
@@
where we define $\triangle^{t}_{2}g(u):=g(u-2t)+g(u+2t)-2g(u)$. Clearly, $t_{k}\to 0$, and letting $a=c^{2}$ (since $a$ is arbitrary), we rewrite (45) as $\triangle_{2}^{t_{k}}g(0)/t_{k}^{2}\to 0$. Apply [36, Chapter 2], Theorem 2.3.1 and conclude: $g^{\prime\prime}(0)=0$, thus $\mathbb{E}[Y^{2}]=0$, and $Y=0$ almost surely, thus $g(u)\equiv 1$. Comparing this with (40), we get: $|f(u)|\equiv 1$. Apply [36, Chapter 2], Theorem 2.1.4, Corollary 1. We get: $f(u)=e^{\mathrm{i}uX_{0}}$ and $X=x_{0}$ almost surely for some $x_{0}\in\mathbb{R}$. But $X$ is $N$-stable. Plugging $X=x_{0}$ into (1), we get: $x_{0}N=cx$. Thus either $N=c$, which contradicts $\mathbb{E}[N]=\infty$; or $x_{0}=0$, which implies $X=0$ almost surely, but we exclude this trivial case. This completes the proof that there is no $N$-stable $X$ in case $\mathbb{E}[N]=\infty$.
@@ 7.2. Proof of Theorem 2
for some $g\in\mathcal{G}$ of index $\alpha$ can be represented as $X=Y^{1/\alpha}Z$, where $Y$ is the standard $N$-stable random variable, and $\mathbb{E}\left[e^{\mathrm{i}uZ}\right]=e^{-g(u)}$. We simply apply [30, Proposition 3.1] and use their remarks on page 309, (10) – (12), Remark 1, about the difference between stable and strictly stable. One can also consult [45, Chapter 1], Section 2 for the latter question.
@@ 7.3. Proof of Theorem 3
If $X$ is a constant, then $M$ and $N$ are also constants, and obviously they commute. If $X$ is not a constant, then $f$ takes at least one value $z$ inside the unit disc $\mathbb{D}=\{z\in\mathbb{C}\mid|z|<1\}$, [36, Chapter 2], Theorem 2.1.4, Corollary 2. But $f$ is continuous, therefore this value $z$ is the limit point of the image of $f$. Both $\varphi\circ\psi$ and $\psi\circ\varphi$ are probability generating functions of some distributions. Therefore, they are analytic on $\mathbb{D}$. Applying the classic result from compelx analysis, we get (16).
@@ 7.7. Proof of Lemma 10
Since $X\geq 0$, by assumption, then the strictly stable random variable $Z$ with index $\alpha$ must be nonnegative as well. But this could be true only if $\alpha=1$ and $Z$ is a positive constant; or $\alpha\in(0,1)$, and $Z$ has Laplace transform (10). This can be found in [51, Chapter V], Theorem 3.5, or in the classic reference [45]. Then the Laplace transform of $X$ is
@@ Acknowledgements
We thank our departmental colleague Tomasz Kozubowski for raising the question whether there exists a Sibuya-stable random variable, and another question whether there exist non-geometric and non-Sibuya commuting semigroups; and further useful discussion. We thank Thierry Huillet for pointing out recent articles on continuous-time branching processes with explicit distributions. We thanks Svetlozar Rachev for multiple useful comments on an earlier draft. We thank Lev Klebanov for useful discussion, and for providing his book [22].